Assume that money demand takes the form 𝑀 𝑃 = π‘Œ[1 βˆ’ (π‘Ÿ + πœ‹ 𝑒 )] where Y = 1000 and r = 0.1.

a. Assume that, in the short run, πœ‹ 𝑒 is constant and equal to 25%. Calculate the amount of seignorage for each rate of money growth, Ξ”M/M, listed below. i. 25% ii. 50% iii. 75%

b. In the medium run, πœ‹ 𝑒 = πœ‹ = Ξ”M/M. Compute the amount of seignorage associated with the three rates of money growth in part (a). Explain why the answers differ from those in part (a).Β 


1
Expert's answer
2021-11-10T11:55:12-0500

(a)In the short run

Signorage = "\Delta" M/M "*M\/P"

If Β is 25% then amount Seigniorage is

M/P= Y(1-(r+Ο€e )

= 1000(1-(0.1+0.25)

=650 (0.25)

= 162.5


If Β is 50% then amount Seigniorage is

=650 (0.50)

=325


If Β is 75% then amount Seigniorage is

=650(0.75)

=487.5


b) In the medium

"\Delta" M/M= Ο€=Ο€e

If Β is 25% then amount Seigniorage is

= 1000(1-(0.1+0.25)

=650 (0.25)

= 162.5


If Β is 50% then amount Seigniorage is

= 1000(1-(0.1+0.5)

=400 (0.5)

=200


If Β is 75% then amount Seigniorage is

= 1000(1-(0.1+0.75)

=150 (0.75)

=112.5

The values differ because overtime, the melt value changes as market demands shift. In the short we assume that Ο€e is fixed while "\Delta" M/M is changing. In the medium run, the two are changing and equal.